OpenStudy (anonymous):

2. Jose is comparing the prices of two carpet cleaning companies. Company A charges $0.40 per square foot of carpet area and an additional $22 per day for services. Company B charges $0.60 per square foot of carpet area and an additional $18 per day for services. Part A: Write equations to represent Company A and Company Bs total charges for cleaning a certain carpet area in a day. Define the variable used in the equations: Part B: Which company would charge less for cleaning 500 square feet of carpet area? Justify your answer. Part C: How much money is saved by using the services of Company A instead of Company B for cleaning 200 square feet of carpet area?

OpenStudy (texaschic101):

PART A : f stands for # of sq ft company A : 0.40f + 22 = total charges per day company B : 0.60f + 18 = total charges per day PART B : company A : .40(500) + 22 = 200 + 22 = 222 for 500 sq ft. company B : .60(500) + 18 = 300 + 18 = 318 for 500 sq ft company A charges less PART C : company A : .40(200) + 22 = 80 + 22 = 102 for 200 sq ft company B : .60(200) + 18 = 120 + 18 = 138 for 200 sq ft Money saved b y using company A : 138 - 102 = 36 any questions ?

OpenStudy (anonymous):

thank you so much

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

you really help you like math

OpenStudy (texaschic101):

I like some math....I hate geometry, but algebra is ok

OpenStudy (anonymous):

texaschic101 can you explain it

OpenStudy (anonymous):

i dont really get the answer and i want to know it too